OpenStudy (anonymous):

What force would you have to apply to a particle which is 3.00 x 10-3 m away from an identical particle with a mass of 5.00 x 10-5 kg?

OpenStudy (anonymous):

F = G m^2/r^2 F is force in Newtons G is universal gravitational constant m is mass in kg r is separation in meters

OpenStudy (anonymous):

For this question, what is G?

OpenStudy (theeric):

I don't get why you're applying a force. But douglaswinslowcooper has an approach to find the gravitational force between them. \(F=G\dfrac{m_1m_2}{r^2}\) \(G\approx 6.674\times 10^{-11}\sf \dfrac{m^3}{s^2\ kg}\) according to http://www.britannica.com/EBchecked/topic/458656/physical-constant#ref204283